### v1

# Network analysis using Krylov subspace trajectories

Abstract
We describe a set of network analysis methods based on the rows of the Krylov subspace matrix computed from a network adjacency matrix via power iteration using a non-random initial vector. We refer to these node-specific row vectors as Krylov subspace trajectories. While power iteration using a random initial starting vector is commonly applied to the network adjacency matrix to compute eigenvector centrality values, this application only uses the final vector generated after numerical convergence. Importantly, use of a random initial vector means that the intermediate results of power iteration are also random and lack a clear interpretation. To the best of our knowledge, use of intermediate power iteration results for network analysis has been limited to techniques that leverage just a single pre-convergence solution, e.g., Power Iteration Clustering. In this paper, we explore methods that apply power iteration with a non-random inital vector to the network adjacency matrix to generate Krylov subspace trajectories for each node. These non-random trajectories provide important information regarding network structure, node importance, and response to perturbations. We have created this short preprint in part to generate feedback from others in the network analysis community who might be aware of similar existing work.
${}^{1}$epartment of Biomedical Data Science
Geisel School of Medicine
Dartmouth College
Hanover, NH 03755, USA
rob.frost@dartmouth.edu
1 Krylov subspace
The order $m$ Krylov subspace [1, 2] for a square diagonalizable $p\times p$ matrix $\mathbf{X}$ and length $p$ vector $\mathbf{v}$ corresponds to the sequence of $m$ vectors generated by the product of $\mathbf{v}$ with powers of $\mathbf{X}$ ranging from 0 to $m-1$, i.e. $\{\mathbf{v},\mathbf{Xv},\mathbf{X^{2}v},...,\mathbf{X^{m-1}v}\}$, and this sequence of vectors can be used to form the columns of a $p\times m$ Krylov subspace matrix $\mathbf{K}$:
$\mathbf{K}=\begin{bmatrix}\vdots&\vdots&\vdots&\vdots&\vdots\\ \mathbf{v}&\mathbf{Xv}&\mathbf{X^{2}v}&\cdots&\mathbf{X^{m-1}v}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \end{bmatrix}$ (1)
We refer to the row vectors of $\mathbf{K}$ as Krylov subspace trajectories with the elements of the trajectory for row $i$ defined by the recursive structure:
$\mathbf{t_{i}}=\{v_{i},\sum_{j=1}^{p}x_{i,j}v_{j},\sum_{j=1}^{p}x_{i,j}(\sum_{k=1}^{p}x_{i,k}v_{k}),...\}$ (2)
For practical numerical applications, the normalized Krylov subspace matrix is used where each column vector is normalized to unit length:
$\mathbf{K}_{n}=\begin{bmatrix}\vdots&\vdots&\vdots&\vdots&\vdots\\ \frac{\mathbf{v}}{||\mathbf{v}||_{2}}&\frac{\mathbf{Xv}}{||\mathbf{Xv}||_{2}}&\frac{\mathbf{X^{2}v}}{||\mathbf{X^{2}v}||_{2}}&\cdots&\frac{\mathbf{X^{m-1}v}}{||\mathbf{X^{m-1}v}||_{2}}\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ \end{bmatrix}$ (3)
In the remainder of this paper, references to the Krylov subspace matrix and associated trajectories will assume that the normalized version is being used.
The Krylov subspace has a number of useful properties and is leveraged in a wide range of numerical linear algebra techniques. In particular, the columns of $\mathbf{K}$ will generally converge to the eigenvector of $\mathbf{X}$ associated with the largest eigenvalue as long as $\mathbf{v}$ is not orthogonal to this principal eigenvector. The columns of $\mathbf{K}$ are also linearly independent so long as $m\leq r+1$ where $r$ is the rank of $\mathbf{X}$. These properties support techniques like Arnoldi iteration [3], which computes the eigenvectors and eigenvalues of $\mathbf{X}$ using the Krylov subspace with incremental orthogonalization. Other Krylov subspace methods include Lanczos iteration [4], the conjugate gradient method [5] and the minimal residual method [6].
2 Krylov subspace and power iteration
The Krylov subspace has a direct correspondence with the incremental results from the method of power iteration [7], whose common realization is defined by Algorithm 1. In particular, the sequence of vectors $\{\mathbf{v}_{0},\mathbf{v}_{1},...,\mathbf{v}_{i}\}$ generated by power iteration correspond to the order $i+1$ normalized Krylov subspace for $\mathbf{X}$ and $\mathbf{v}_{0}$. Note that the specification of power iteration in Algorithm 1 does not include features important for many numerical applications, e.g, checks to ensure that $\mathbf{X}_{i}$ is well conditioned, stochastic initialization of $\mathbf{v}_{0}$, use of methods like accelerated stochastic power iteration [8] to improve computational performance, alternate stopping conditions, etc..
Algorithm 1 Power iteration
Input:
Output:
1:for $i\in\{1,...,maxIter\}$ do
2:     $\mathbf{v}_{i}=\mathbf{X}\mathbf{v}_{i-1}$triangleright Update $\mathbf{v}_{i}$
3:     $\mathbf{v}_{i}=\mathbf{v}_{i}/||\mathbf{v}_{i}||_{2}$triangleright Normalize $\mathbf{v}_{i}$ to unit length
4:     $\lambda_{i}=\mathbf{v}_{i}^{T}\mathbf{X}\mathbf{v}_{i}$triangleright Compute principal eigenvalue estimate
5:     if $i>1$ then
6:         $\delta=(|\lambda_{i-1}-\lambda_{i}|)/\lambda_{i}$triangleright Compute proportional change in eigenvalue estimate
7:         if $\delta<tol$ then
8:              break $\triangleright$ If proportion change is less than $tol$, exit               return $\mathbf{v}_{i},\lambda_{i},i$
3 Existing uses of the Krylov subspace for network analysis
In the network analysis space, the Krylov subspace is most commonly generated through the application of power iteration to the network adjacency matrix, $\mathbf{A}$, which, for an edge-weighted network defined over $p$ nodes is the $p\times p$ matrix:
$\mathbf{A}=\begin{bmatrix}a_{1,1}&\cdots&a_{1,p}\\ \vdots&\ddots&\vdots\\ a_{p,1}&\cdots&a_{p,p}\end{bmatrix}$ (4)
where $a_{i,j}$ holds the weight of the edge between nodes $i$ and $j$ or 0 if no such edge exists. Self-edges are represented by the diagonal elements. If the network is undirected, $a_{i,j}=a_{j,i}$ and $\mathbf{A}$ is symmetric; if the network is directed, then $a_{i,j}$ and $a_{j,i}$ capture distinct edges and $\mathbf{A}$ is asymmetric.
Power iteration is normally applied to the network adjacency matrix using a random $\mathbf{v}_{0}$ to compute eigenvector centrality values (or values for a related centrality measure) [9]. Eigenvector centrality models node importance as the weighted sum of the importance of adjacent nodes and therefore leads to the following eigenvalue problem:
$\mathbf{A}\mathbf{v}=\lambda\mathbf{v}$ (5)
Specifically, the eigenvector centrality for node $n$ is equal to element $n$ of the eigenvector $\mathbf{v}$ corresponding to the largest eigenvalue. When the network is strongly connected the matrix $\mathbf{A}$ is irreducible and the Perron-Frobenius theorem [10] guarantees that there is a unique largest real eigenvalue with a corresponding eigenvector that has strictly positive elements.
While this application of power iteration generates the $k+1$ order Krylov subspace of $\mathbf{A}$, where $k$ is the number of iterations until numerical convergence, computation of eigenvector centrality only uses the final vector. In addition, the fact that a random $\mathbf{v}_{0}$ is almost always employed means the generated trajectories are also random and therefore lack a clear interpretation. Other network analysis techniques based on the Krylov subspace of $\mathbf{A}$ (or a related matrix) also use just a single intermediate solution. These techniques include the propagation of node information in Graph Convolution networks via a single iteration of power iteration [11] and the Power Iteration Clustering method [12] that performs a truncated power iteration using a random initial vector on a normalized affinity matrix and then uses the last eigenvector estimate to perform node clustering via the k-means algorithm.
4 Krylov subspace trajectories of the network adjacency matrix
In this paper, we are interested in the Krylov subspace associated with a network adjacency matrix $\mathbf{A}$ and non-random initial vector $\mathbf{v}_{0}$. Generation of this subspace using the fixed initial vector $\mathbf{v}_{0}=\{1/\sqrt{p},...,1/\sqrt{p}\}$ is specified by Algorithm 2.
Algorithm 2 Krylov subspace of network adjacency matrix
Input:
Output:
Notation:
$\mathbf{K}\in\mathbb{R}^{p\times maxIter+1}$triangleright Initialize $\mathbf{K}$ as a $p\times maxIter+1$ matrix
[…]
3:for $i\in\{1,...,maxIter\}$ do
4:     $\mathbf{K}[,i+1]=\mathbf{X}\mathbf{K}[,i]$triangleright Compute $i+1$ column of $\mathbf{K}$
5:     $\mathbf{K}[,i+1]=\mathbf{K}[,i+1]/||\mathbf{K}[,i+1]||_{2}$triangleright Normalize column to unit length
6:     $\lambda_{i}=\mathbf{K}[,i+1]^{T}\mathbf{X}\mathbf{K}[,i+1]$triangleright Compute principal eigenvalue estimate
7:     if $i>1$ then
8:         $\delta=(|\lambda_{i-1}-\lambda_{i}|)/\lambda_{i}$triangleright Compute proportional change in eigenvalue estimate
9:         if $\delta<tol$ then
10:              break $\triangleright$ If proportion change is less than $tol$, exit               return $\mathbf{K}[,\{1,...,i+1\}]$
To illustrate an example of a non-random network Krylov subspace, Algorithm 2 was applied to an unweighted and undirected tree network (visualized in Figure 1) that was generated with a total of 31 nodes and two child nodes per non-leaf node. For this network, Algorithm 2 converged to an estimated eigenvalue of 2.4 after 16 iterations. The order 17 Krylov subspace trajectories of the 31 nodes in this network are visualized in Figure 2. An interesting feature of the trajectories for this tree graph is the presence of oscillations. For this particular network, the power iteration method (as defined in Algorithms 1 and 2) with a uniform starting vector does not in fact converge to the largest eigenvector/eigenvalue pair but instead halts at an eigenvalue estimate of 2.4 (the true largest eigenvalue is 2.4495) with alternating eigenvector estimates $\mathbf{v}_{a}$ and $\mathbf{v}_{b}$, i.e., $\mathbf{A}\mathbf{v}_{a}=\alpha\mathbf{v}_{b}$, $\mathbf{A}\mathbf{v}_{b}=\beta\mathbf{v}_{a}$, and $\mathbf{v}_{a}^{T}\mathbf{A}\mathbf{v}_{a}=\mathbf{v}_{b}^{T}\mathbf{A}\mathbf{v}_{b}=2.4$). For the more complex edge-weighted networks associated with experimental data, power iteration is unlikely to encounter such erroneous convergence behavior. Although this behavior makes a simple implementation of power iteration ineffective for eigenvector centrality computation on this network, the generated Krylov subspace trajectories still contain important information about network structure. The utility of these trajectories is explored in more detail in Sections 5-7.
Figure 1: Example tree network. The left panel visualizes the network with equal sized nodes. The right panel visualizes with node size based on eigenvector centrality.
Figure 2: Node-specific Krylov trajectories for the tree network visualized in Figure 1.
5 Network clustering using Krylov subspace trajectories
One of the key applications of the Krylov subspace trajectories generated by Algorithm 2 is node clustering or community detection. Importantly, clustering of nodes based on these trajectories yields very distinct results from standard approaches for network community detection such as Louvain clustering [13]. To illustrate this phenomenon, the nodes in tree network visualized in Figure 1 were clustered using both Louvain clustering (as implemented by the igraph R package [14] function cluster_louvain()) and via hierarchical agglomerative clustering using correlation distance between the Krylov subspace trajectories and complete linkage (the dendrogram was cut to yield four clusters). These two cluster solutions are visualizied in Figure 3. For this example, Louvain clustering generates an expected solution that groups nodes according to relative proximity in the graph. Clustering using the Krylov subspace trajectories, on the other hand, generates a very distinct solution that groups nodes based on similar structural position in the network rather than proximity. While the revealed structural-based groups are obvious for the simple tree network, this type of analysis can yield important non-obvious node relationships in the more complex networks used to model real world data.
Figure 3: Clustering of the example tree network generated by Louvain clustering (left panel) and hierarchical agglomerative clustering using correlation distance between the Krylov subspace trajectories and complete linkage (right panel).
6 Perturbation analysis using Krylov subspace trajectories
Algorithm 2 uses the uniform initial vector $\mathbf{v}_{0}=\{1/\sqrt{p},...,1/\sqrt{p}\}$ to compute the Krylov subspace trajectories, which is motivated for applications that treat each node equally. For some applications, however, unequal weights can be associated with network nodes and these node weights can be used to create a non-uniform (though still non-random) $\mathbf{v}_{0}$. One example of such an application would be modeling of a perturbation applied to specific network nodes. To illustrate this type of perturbation analysis, Algorithm 2 was applied to the tree network using a $\mathbf{v}_{0}$ where the elements for nodes 5 and 6 are five times larger than the elements for other nodes. The trajectories generated in this case are visualized in Figure 4. Clustering based on these perturbed trajectories (shown in Figure 5) yields a distinct grouping of nodes that capture both overall structural position and response to the node 5 and 6 perturbations.
Figure 4: Node-specific Krylov trajectories for the tree network visualized in Figure 1 using a non-uniform $\mathbf{v}_{0}$ where the entries for nodes 5 and 6 that are five times larger than the entries for other nodes.
Figure 5: Clustering of the example tree network generated by hierarchical agglomerative clustering using correlation distance between the Krylov subspace trajectories and complete linkage. The left panel visualizes the clusters when $\mathbf{v}_{0}$ is uniform (trajectories are shown in Figure 2) and the right panel visualizes the clusters when $\mathbf{v}_{0}$ has inflated entries for nodes 5 and 6 and that five times larger than the values for other nodes (trajectories are shown in Figure 4).
7 Capturing node importance using Krylov subspace trajectories
Krylov subspace trajectories can also be used to capture node importance. One approach to this involves the computation of a single summary statistic for each trajectory. Although many possible summary statistics exist, we believe statistics that quantify the level of oscillations in the trajectory may have utility. One simple way to quantify this is to subtract the absolute difference between the starting and ending values of a given trajectory from the sum of the absolute values of all sequential differences; we refer this trajectory statistic $\delta$. The $\delta$ statistic captures the number and magnitude of oscillations in the trajectory (for monotonic trajectories, it is 0). If the vector $\mathbf{t}$ captures the length $i+1$ trajectory for a given node, then $\delta$ is computed as:
$\delta=(\sum_{j=2}^{j=i+1}|t_{j}-t_{j-1}|)-|t_{i+1}-t_{1}|$ (6)
Figure 6 visualizes the $\delta$ statistic for the example tree network as computed using both uniform and non-uniform $\mathbf{v}_{0}$.
Figure 6: Visualization of the $\delta$ statistic for the tree network as computed on the Krylov subspace trajectories for uniform $\mathbf{v}_{0}$ (left panel; trajectories shown in Figure 2) or non-uniform $\mathbf{v}_{0}$ (right panel; trajectories shown in Figure 4). To enable visualization for both cases, the $\delta$ values are divided by the maximum $\delta$ for that network.
Acknowledgments
This work was funded by National Institutes of Health grants R35GM146586, R21CA253408, P20GM130454 and P30CA023108. We would like to acknowledge the supportive environment at the Geisel School of Medicine at Dartmouth where this research was performed.